Fermi National Laboratory
 

Accelerator Update for November 14, 2007 - November 16, 2007

 

Wednesday November 14

The day shift began with the Tevatron (TeV) conditioning its separators, with Operations monitoring the Recycler stash (198.1E10), with the Antiproton Source (Pbar) stacking (259.5E10), with the Main Injector (MI), Booster, and Linac operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for beam, and with the SciBooNE / MiniBooNE experiments taking beam.

Operations prepared Pbar for the first of several accesses throughout the day by the T972 experiment. (The T972 experiment is testing the effectiveness of shielding.)

At 10:26 AM, the NuMI Horn LCW system alarmed on a high level alarm. Operations contacted an expert who said they were working on the system.

At 10:47 AM, the750 KeV area buncher tripped off. It reset.

An LCW expert reported at 11:10 AM that he was valving in column three at CUB.

A MI expert began a slipstacking study at 1:43 PM.

A TeV expert took the separators out of conditioning mode at 2:08 PM.

Operations established a proton only store at 3:12 PM.

Operations put the TeV into shot setup at 3:46 PM.

At 5:20 PM, a Linac RF station (LRF2) and the Linac buncher tripped off. They both reset.

At 7:02 PM, a TeV separator sparked at sector C49. This caused a small luminosity hit in the upcoming store.

Operations established store 5722 at 7:05 PM with an initial luminosity of 103.37E30.

At 7:45 PM, LRF5 tripped off for the first of two times in 70 minutes. It reset.

Operations reported at 11:57 PM that the Recycler had stashed antiprotons once during the evening shift.

Thursday November 15

The midnight shift began with Operations monitoring store 5722 (57.9E30) and the stash (153.1E10), with Pbar stacking (182.3E10), and with Sci /MiniBooNE taking beam.

At 12:36 AM, a MI RF station (MIRF1) tripped off and would not reset. Operators bypassed the station.

At 1:32 AM, the 750 KeV area buncher tripped off. It reset.

Operations terminated store 5722 at 3:01 AM.

Operations prepared D0 for an access at 3:17 AM. They came out of access at 5:49 AM.

Operations put the TeV through a dry squeeze at 6:10 AM.

Operations put the TeV into shot setup at 6:38 AM.

A Linac expert began a study period at 8 AM. He reported that LRF5 would be off for this study.

Operations began preparing MI for an access at 8:12 AM to repair LCW leaks on the bus around the collimators. Experts reported that this work might take 12 hours.

Operations established store 5724 at 8:26 AM with an initial luminosity of 112.3E30.

At 8:43 AM, the 750 KeV area buncher tripped off. It reset.

At 8:48 AM, LRF4 tripped off. It reset.

At 9:25 AM, LRF3 & 4 tripped off. Both stations reset.

At 9:46 AM, the Linac buncher tripped off. It reset.

Operations began preparing Sci/MiniBooNE for access at 9:54 AM.

The Neutron Therapy Facility reported at 11:10 AM that they had completed treatment on their last patient for the day.

Sci/MiniBooNE came out of access at 11:24 AM.

At 11:41 AM, BRF17 tripped off for the first of three times in 12 minutes. It reset.

A Linac expert completed his tuning and Operations began sending Booster beam to its absorber at 12:08 PM.

Operations resumed sending beam to Sci/MiniBooNE at 4:09 PM.

At 4:55 PM, a Linac Klystron RF station (KRF2) tripped off on an overcurrent indication. Experts investigated.

At 5:55 PM, Operations reported problems with the Booster vacuum alarms. A controls expert investigated and rebooted a process.

MI came out of access at 7:21 PM, but Operations could not get a beam permit due to a ground fault indication.

Operations reported at 8:10 PM that an expert replaced a Booster power supply for VBCO.

At 8:20 PM, a Linac expert reported that the KRF2 charging power supply transformer would need to be replaced.

Operations reported at 9:15 PM that the MI ground fault was found; a loose cable touching a bus near the upstream collimator caused the fault.

A Linac expert reported at 10:46 PM that he would replace a power tube for the 750 KeV area buncher while others were working on the KRF2 charging supply.

Friday November 16

The midnight shift began with Operations monitoring store 5724 (29.19E30) and a stack (234.4E10), and with experts working on KRF2.

A Linac expert reported at 3:11 AM that they had completed work on KRF2, but they found a LCW leak on the back of the supply.

Operations established beam to the Linac absorber at 4:23 AM.

Operations resumed sending beam to Sci/MiniBooNE at 4:25 AM.

Operation began sending beam to the Recycler at 4:32 AM.

At 5:30 AM, LRF5 tripped off. It reset.

At 6:10 AM, Operations reported an indeterminate status on a Recycler beam valve. It would show closed, but not open. Operations had stashed the Recycler just 30 minutes ago. An expert investigated to see if the problem was due to a bad control card (in the service building) or due to a bad solenoid or microswitch (in the tunnel).

At 7:49 AM, the Recycler vacuum expert reported that he had bypassed the controller, but could not force the valve to open; it goes from closed to indeterminate status.

The Plans for Friday and the Weekend

The plans for today are to fix the Recycler beam valve problem. The Run Coordinator wants a store in the TeV by this afternoon. The plans for the weekend are to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/16/2007 -->    email Fermilab